OpenStudy (anonymous):

i need help with a story problem. Peyton has 6 times as many pencils and paintbrushes. He has a total of 63 pencils and paintbrushes. How many of each does Peyton have? Thanks

OpenStudy (anonymous):

9 and 54

OpenStudy (anonymous):

You can use a system of equations. Say p is pencils and t is paintbrushes(the number of) Then you have p+t=63 and p=6t Plug in p to get 6t+t=63 7t=63 t=9. He has 9 paint brushes and 6 times that number of pencils so he has 54 pencils

OpenStudy (anonymous):

x for paintbrush 6x for pencil

OpenStudy (anonymous):

since the total is 63, x + 6x = 63

OpenStudy (anonymous):

7x = 63

OpenStudy (anonymous):

x = 9

OpenStudy (anonymous):

and 6x = 6(9) = 54